OpenStudy (anonymous):

\[Integrate = \int\limits_{e}^{e ^{3}} \frac{dx}{x}\]

OpenStudy (lgbasallote):

first step...what is the integral of dx/x?

OpenStudy (mimi_x3):

\[\int\limits \frac{1}{x} dx => \ln(x) +C\]its a standard integral.. then sub in the limits.

OpenStudy (anonymous):

what is the \[\ln e\]??? and \[\ln e ^{3}\]???

OpenStudy (mimi_x3):

\[\large \left[\ln(x)\right] _{e}^{e^{2}} => \ln(e^{2} )- \ln(e)\] you can use log laws to simplify it.. howeber, \(ln(e)=1\)

OpenStudy (anonymous):

ok thanks @Mimi_x3 i got the answer. :)

OpenStudy (mimi_x3):

np